Bill & Jose's Discount Cinema sold adults' tickets for $4.10 each and children's tickets for $2.70 each. Last Tuesday, a total of $331.30 was collected from 89 movie watchers. How many of each type of ticket were sold? (Let x represent the number of adults' tickets and y represent the number of children's tickets.)
Answer by stanbon(75887)

Bill & Jose's Discount Cinema sold adults' tickets for $4.10 each and children's tickets for $2.70 each. Last Tuesday, a total of $331.30 was collected from 89 movie watchers. How many of each type of ticket were sold? (Let x represent the number of adults' tickets and y represent the number of children's tickets.)
----------------
Equations:
Quantity: x + y = 89
Value ::4.1x+2.7y = 331.30
--------------------------------
Multiply thru 1st eq. by 410.
Multiply thru 2nd eq. by 100 to get:
------------------------------------------
410x + 410y = 410*89
410x + 270y = 33130
----------------------------
Subtract 2nd eq from 1st and solve for "y":
40y = 3360
y = 84 (# of children tickets sold)
---
Since x + y = 89, x = 5 (# of adult tickets sold)
=====================
Cheers,
Stan H.